OpenStudy (anonymous):

Solve the system of equations by substitution method (-5x-15y=9) (5x+15y=0)

hero (hero):

Rewrite the first equation: 5x + 15y = -9 So now you have: 5x + 15y = -9 5x + 15y = 0 So you know both lines are parallel, therefore, it is not possible to solve the system using substitution method nor any other method.

OpenStudy (anonymous):

Thank you so much.